# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ 2.2 The Model Capacity and Task Difficulty
Suppose $\lambda_{j}\eqsim j^{-\beta}$ for some $\beta\in(1,\infty)$.
This condition ensures $\operatorname{Tr}(\mathbf{H}):=\sum_{j=1}^{N}\lambda_{j}<\infty$ and thus, the RKHS $\mathcal{H}_{k}$ is well defined. A smaller $\beta$ implies slower eigenvalue decay, hence a larger RKHS $\mathcal{H}_{k}$. Consequently, the expressiveness of the student model (1) is governed by two orthogonal factors: model size $M$ (the number of features used) and capacity exponent $\beta$ (the intrinsic richness of the features). Conceptually, $\beta$ can be viewed as an architecture-dependent quantity: models with different architectures may differ substantially in expressiveness even they have the same number of parameters.
@@
Suppose $\theta^{*}_{j}\eqsim j^{-1/2}\lambda_{j}^{(s-1)/2}$ for all $j\in[N]$ for some $s\in(0,\infty)$.
Under Assumptions 2.2 and 2.3, the target function can be expressed as
$f^{*}=\sum_{j=1}^{N}\theta_{j}^{*}\phi_{j}\eqsim\sum_{j=1}^{N}j^{-1/2}\lambda_{j}^{s/2}e_{j}\stackrel{{\scriptstyle(i)}}{{\eqsim}}\sum_{j=1}^{N}j^{-(s\beta+1)/2}e_{j},$
where $(i)$ follows from the assumption $\lambda_{j}\eqsim j^{-\beta}$. Thus, the parameter $\alpha:=s\beta$ characterizes the task’s intrinsic difficulty, while $s$ measures the relative difficulty with respect to the chosen model. For a fixed target (i.e., fixed $\alpha$), choosing a higher-capacity model (smaller $\beta$) increases the value of $s=\alpha/\beta$, effectively making the task relatively easier. We will analyze, for a fixed target, how varying $\beta$, i.e., the model capacity, influences the corresponding scaling behavior.
@@
We remark that similar capacity and source conditions have been widely used in the analysis of kernel methods (caponnetto2005fast, ; caponnetto2007optimal, ; spigler2020asymptotic, ; bordelon2020spectrum, ; maloney2022solvable, ), and recently applied to study their scaling behavior (paquette2024fourplus, ; bordelon2024feature, ). Our work builds upon and extends this line of research.
2.3 Online Minibatch SGD
@@
Given a data point $\mathbf{z}=(\mathbf{x},y)\in\mathbb{R}^{d}\times\mathbb{R}$ and a student model $f(\cdot;\mathbf{v})$, we use the square loss to measure fitting error $\ell(\mathbf{z},\mathbf{v}):=\frac{1}{2}(f(\mathbf{x};\mathbf{v})-y)^{2}$. Then, the population risk is given by
$\mathcal{R}(\mathbf{v}):=\mathbb{E}_{\mathbf{z}}[\ell(\mathbf{z},\mathbf{v})]=\frac{1}{2}\|\mathbf{W}^{\top}\mathbf{v}-\bm{\theta}^{*}\|_{\mathbf{H}}^{2}+\frac{\sigma^{2}}{2}=:\mathcal{E}(\mathbf{v})+\frac{\sigma^{2}}{2},$ (2)
@@
where $\mathcal{E}(\mathbf{v})$ is the excess risk.
We consider to minimize $\mathcal{R}(\cdot)$ using online minibatch SGD. Specifically, at each step, we randomly sample a batch of i.i.d. samples $S_{k}:=\{(\mathbf{x}_{k,j},y_{k,j})\}_{j=1}^{B_{k}}$, where $B_{k}$ is the batch size at the $k$-th step. Then, SGD iterates as follows:
$\mathbf{v}_{k}:=\mathbf{v}_{k-1}-\frac{\eta_{k}}{B_{k}}\sum\nolimits_{\mathbf{z}\in S_{k}}\nabla_{\mathbf{v}}\ell(\mathbf{z},\mathbf{v}_{k-1}),$ (3)
where $\mathbf{v}_{k}$ denotes the weight of student model at the $k$-th step with $\mathbf{v}_{0}=\bm{0}$, and $\eta_{k}\geqslant 0$ is the learning rate at the $k$-th step. Rewrite the SGD update (3) as:
$\mathbf{v}_{k}=\mathbf{v}_{k-1}-\eta_{k}\left(\nabla\mathcal{R}(\mathbf{v}_{k-1})+\bm{\xi}_{k}\right)$ (4)
@@
where $\bm{\xi}_{k}:=\frac{1}{B_{k}}\sum_{\mathbf{z}\in S_{k}}\nabla\ell(\mathbf{z},\mathbf{v}_{k-1})-\nabla\mathcal{R}(\mathbf{v}_{k-1})$ denotes the gradient noise that satisfies $\mathbb{E}[\bm{\xi}_{k}]=0,\mathbb{E}[\bm{\xi}_{k}\bm{\xi}_{k}^{\top}]=\frac{1}{B_{k}}\Sigma(\mathbf{v}_{k-1}),$ where $\Sigma(\mathbf{v})$ denotes the noise covariance at $\mathbf{v}$ for the batch size $1$.
Throughout this paper, we refer to $\bm{\eta}:=(\eta_{1},\eta_{2},\cdots,\eta_{K})^{\top}\in\mathbb{R}_{\geqslant 0}^{K}$ as the learning rate schedule (LRS). In practical training of large-scale deep models, popular LRSs include the cosine schedule (loshchilov2016sgdr, ; touvron2023llama, ), the warmup-stable-decay (WSD) schedule (hu2024minicpm, ), and the multistep schedule (bi2024deepseek, ) (see details in Appendix B.1). In addition, we refer to $\mathbf{b}=(B_{1},B_{2},\dots,B_{K})^{\top}\in\mathbb{N}_{+}^{K}$ as the batch size schedule.
A key insight for tractability is that the gradient noise exhibits the following anisotropic structure, whose proof is deferred to Appendix C.1.

### v3

@@ 2.1 Model Capacity and Task Difficulty
The spectrum of the feature map satisfies $\lambda_{j}\eqsim j^{-\beta},\beta>1.$
The condition $\beta>1$ ensures $\mathrm{tr}(\mathbf{H})=\sum_{j=1}^{N}\lambda_{j}\leqslant C$ for some constant $C$ independent of $N$, making our analysis dimension-free and applicable to the infinite-dimensional setting ($N=\infty$).
Restricting attention to the top-$M$ features, any model of the form
$f(\cdot;\mathbf{v})=\sum_{j=1}^{M}v_{j}\phi_{j}=\sum_{j=1}^{M}v_{j}\lambda_{j}^{1/2}\,\widehat{\phi}_{j}\eqsim\sum_{j=1}^{M}v_{j}\,j^{-\beta/2}\widehat{\phi}_{j}$
reveals that higher-index features are increasingly down-weighted by the factor $j^{-\beta/2}$. As $\beta$ increases, the spectrum decays more rapidly, and the model effectively relies on fewer features. Hence, the model’s expressive power is governed by two complementary factors:
  • the model size $M$, which controls how many features are retained, and
  • the capacity exponent $\beta$, which controls how quickly these features decay in importance.
Remark 2.4.
Note that for a fixed target function $f^{*}$, one can adopt different (potentially) non-linear feature map $\bm{\phi}$ (and hence different $\beta$). The value of $\beta$ reflects the capacity of the chosen features. For instance, consider $\bm{\phi}(\mathbf{x})=\nabla_{\theta}\mathcal{N}(\mathbf{x};\theta)$, where $\mathcal{N}(\cdot;\theta)$ denotes a neural network. Then $\bm{\phi}(\mathbf{x})$ corresponds to neural tangent features, and the associated kernel $K_{\phi}(\mathbf{x},\mathbf{x}^{\prime}):=\bm{\phi}(\mathbf{x})^{\top}\bm{\phi}(\mathbf{x}^{\prime})$ is known as the neural tangent kernel (NTK) [24]. In this case, the network depth and activation functions govern the spectral decay, and thus determine the effective exponent $\beta$ [6, 67].
@@
Suppose $|\theta^{*}_{j}|^{2}\eqsim j^{-1}\lambda_{j}^{\,s-1}$ for some $s>0$.
Under Assumptions 2.3 and 2.5, the target function admits the expansion
$f^{*}=\sum_{j=1}^{N}\theta_{j}^{*}\phi_{j}\eqsim\sum_{j=1}^{N}j^{-1/2}\lambda_{j}^{\,s/2}\;\widehat{\phi}_{j}\eqsim\sum_{j=1}^{N}j^{-(s\beta+1)/2}\;\widehat{\phi}_{j}.$
Since $\{\hat{\phi}_{j}\}$ are orthonormal, this assumption implies that the spectral energy of $f^{*}$ decays as a power law. The exponent $\alpha:=s\beta$ therefore quantifies the task’s intrinsic difficulty, which depends only on the target function itself and is independent of the model spectrum. In contrast, $s$ measures the relative difficulty with respect to a model of capacity $\beta$: for a fixed $f^{*}$ (fixed $\alpha$), adopting a higher-capacity model (smaller $\beta$) increases $s=\alpha/\beta$, making the task relatively easier. In other words, the same task appears easier to a higher-capacity model.
@@
We remark that similar assumptions—commonly referred to as capacity and source conditions—have been widely used in the analysis of kernel methods [13, 12, 58, 10, 41]. Our work builds upon and extends this line of research.
3 One-Pass SGD and Intrinsic-Time SDE
@@
Given a data point $\mathbf{z}=(\mathbf{x},y)\in\mathcal{X}\times\mathbb{R}$ and a model $f(\cdot;\mathbf{v})$, define the loss $\ell(\mathbf{z},\mathbf{v})=\tfrac{1}{2}\bigl(f(\mathbf{x};\mathbf{v})-y\bigr)^{2}.$ Then, the population risk is
$\mathcal{R}(\mathbf{v})=\mathbb{E}_{\mathbf{z}}[\ell(\mathbf{z},\mathbf{v})]=\tfrac{1}{2}\|\mathbf{W}^{\top}\mathbf{v}-\bm{\theta}^{*}\|_{\mathbf{H}}^{2}+\tfrac{\sigma^{2}}{2}=:\mathcal{E}(\mathbf{v})+\tfrac{\sigma^{2}}{2},$ (3)
@@
where $\mathcal{E}(\mathbf{v})$ denotes the excess risk. We minimize $\mathcal{R}(\mathbf{v})$ via one-pass SGD, given by
$\mathbf{v}_{k+1}=\mathbf{v}_{k}-\frac{\eta_{k}}{B_{k}}\sum_{\mathbf{z}\in S_{k}}\nabla_{\mathbf{v}}\ell(\mathbf{z},\mathbf{v}_{k}),$ (4)
where $S_{k}:=\{(\mathbf{x}_{k,j},y_{k,j})\}_{j=1}^{B_{k}}$ is a mini-batch of i.i.d. samples, $\eta_{k}$ and $B_{k}$ are the learning rate and batch size, respectively. The initialization is set to $\mathbf{v}_{0}=\bm{0}$.
Throughout, we refer to $\bm{\eta}:=(\eta_{0},\eta_{1},\ldots,\eta_{K-1})$ as the learning rate schedule (LRS). Common choices in practice include the cosine [39, 61], WSD [22], and multi-step [5] schedules (see Appendix A.2 for details). To analyze the effect of LRS, we rewrite (4) as
$\mathbf{v}_{k+1}=\mathbf{v}_{k}-\eta_{k}\bigl(\nabla\mathcal{R}(\mathbf{v}_{k})+\bm{\xi}_{k}\bigr),$ (5)
@@
where the gradient noise $\bm{\xi}_{k}=\tfrac{1}{B_{k}}\sum_{\mathbf{z}\in S_{k}}\nabla\ell(\mathbf{z},\mathbf{v}_{k})-\nabla\mathcal{R}(\mathbf{v}_{k})$ satisfies $\mathbb{E}[\bm{\xi}_{k}]=0,\mathbb{E}[\bm{\xi}_{k}\bm{\xi}_{k}^{\top}]=\frac{1}{B_{k}}\bm{\Sigma}(\mathbf{v}_{k}),$ with $\bm{\Sigma}(\cdot)$ denoting the noise covariance for batch size $1$.
Remark 3.1.
The hypercontractivity condition on the feature map $\bm{\phi}$ (Assumption 2.1) ensures that the noise covariance $\bm{\Sigma}(\cdot)$ can be characterized up to multiplicative constants by an explicit analytical expression; see Lemma 4.8.
Continuous-time limit.
Following prior work [32, 33, 34, 35, 52, 36], we analyze the continuous-time limit of SGD rather than the discrete update (4) or (5). This perspective makes the analysis more tractable and clarifies the emergence of scaling laws. Fix a discretization step size $h>0$ and let $\varphi_{k}:=\eta_{k}/h$ for $k\in\mathbb{N}$. Then, (5) becomes $\mathbf{v}_{k+1}=\mathbf{v}_{k}-\varphi_{k}\nabla\mathcal{R}(\mathbf{v}_{k})h-\varphi_{k}h\bm{\xi}_{k}.$ For sufficiently small $h$, this iteration is well approximated by the Itô-type SDE [33, 47]:
$\mathop{}\!\mathrm{d}\bar{\mathbf{v}}_{\tau}=-\varphi(\tau)\nabla\mathcal{R}(\bar{\mathbf{v}}_{\tau})\mathop{}\!\mathrm{d}\tau+\varphi(\tau)\sqrt{\frac{h}{b(\tau)}\bm{\Sigma}(\bar{\mathbf{v}}_{\tau})}\mathop{}\!\mathrm{d}\mathbf{B}_{\tau},$ (6)
where $\mathbf{B}_{\tau}\in\mathbb{R}^{M}$ is an $M$-dimensional Brownian motion, and
  • $\varphi(\cdot)$ is the continuous-time LRS satisfying $\varphi(kh)=\eta_{k}/h$ for all $k\in\mathbb{N}$;
  • $b(\cdot)$ is the continuous-time batch-size schedule satisfying $b(kh)=B_{k}$ for all $k\in\mathbb{N}$.
In (6), the learning rate affects both the drift and diffusion terms, thereby coupling the deterministic and stochastic effects.
Intrinsic-time reparametrization.
In SDE (6), the physical time $\tau$ serves as the continuous analogue of the discrete step index $k$. However, when the learning rate varies over time, the actual training progress is determined not by the number of updates $k$ but by the accumulated step size $\sum_{j=1}^{k}\eta_{j}$, which more faithfully reflects the total optimization effort. Motivated by this observation, we introduce an intrinsic time variable that rescales the physical time $\tau$ according to the LRS:
$t=T(\tau):=\int_{0}^{\tau}\varphi(r)\mathop{}\!\mathrm{d}r,$ (7)
which measures the LRS-adjusted training duration. Let $\bm{\nu}_{t}=\bar{\mathbf{v}}_{T^{-1}(t)}.$ Applying the Øksendal’s time change formula [46] to the SDE (6) yields the following intrinsic-time SDE:
$\mathop{}\!\mathrm{d}\bm{\nu}_{t}=-\nabla\mathcal{R}(\bm{\nu}_{t})\mathop{}\!\mathrm{d}t+\sqrt{\gamma(t)\,\bm{\Sigma}(\bm{\nu}_{t})}\mathop{}\!\mathrm{d}\mathbf{B}_{t}\,\,\text{ with }\,\,\gamma(t)=\frac{h\varphi(T^{-1}(t))}{b(T^{-1}(t))}.$ (8)
Here $\gamma(t)$ quantifies the joint effect of learning-rate and batch-size scheduling. Compared with (6), the LRS dependence is absorbed from the drift and retained only in the diffusion term, thereby decoupling the deterministic and stochastic effects. This structural simplification greatly facilitates the subsequent scaling analysis.
For a clearer explanation of the connection between the discrete SGD (5) and the SDE formulations (6) and (8), we refer the reader to Appendix A.4.
4 Intrinsic-Time Functional Scaling Laws
In this section, we present our main results on functional scaling laws. A high-level outline of the core proof ideas is provided in Section 4.4, while all technical details are deferred to Appendix C.
We begin with assumptions on the learning-rate schedule and model size.
Assumption 4.1.
Assume that Assumptions 2.1, 2.3, and 2.5 hold, $M$ is sufficiently large with $N-M\gtrsim M$, and the LRS satisfies $\sup_{t\geqslant 0}\gamma(t)\leqslant C$ for some sufficiently small constant $C>0$.
4.1 An Illustrative Setting and the Underlying Scaling Behavior
Theorem 4.2 (Intrinsic-Time FSL, hard-regime).
Under Assumption 4.1, let $\bm{\nu}_{t}$ denote the solution to the intrinsic-time SDE (8) with top-$M$ features. Then, for any $f^{*}$ with difficulty $s\in(0,1-1/\beta]$ and any $\sigma\geqslant 0$, it holds for all $t\geqslant 0$ that $\mathbb{E}[\mathcal{R}(\bm{\nu}_{t})]-\tfrac{1}{2}\sigma^{2}\eqsim M^{-s\beta}+e(t)+\int_{0}^{t}\mathcal{K}(t-z)[e(z)+\sigma^{2}]\gamma(z)\mathop{}\!\mathrm{d}z,$ (9) where $e(t):=(1+t)^{-s},\;\mathcal{K}(t):=(1+t)^{-(2-1/\beta)}$. For the random-$M$ case, the same FSL holds with probability at least $1-\exp\{-\Omega(M)\}$ over the randomness of $\mathbf{W}$.
This theorem establishes that, for hard tasks with $s\leqslant 1-1/\beta$, the loss dynamics are fully characterized by the FSL (9). Moreover, each term in the FSL (9) admits a clear interpretation:
  • Irreducible error: $\tfrac{1}{2}\sigma^{2}$. This term is due to label noise.
  • Approximation error: $M^{-s\beta}$. This term corresponds to the error due to finite model size, with the scaling efficiency is determined by the task’s intrinsic difficulty $s\beta$.
  • Signal learning: $e(t)$. This term corresponds to learning under full-batch gradient descent, capturing the rate at which SGD extracts the signal $f^{*}$. Moreover, the rate depends on the task’s relative difficulty $s$. For a fixed target $f^{*}$ (fixed $\alpha=s\beta$), increasing model capacity (smaller $\beta$) accelerates its convergence since $s=\alpha/\beta$ becomes larger.
  • Noise accumulation: $\int_{0}^{t}\mathcal{K}(t-z)[e(z)+\sigma^{2}]\gamma(z)\mathop{}\!\mathrm{d}z$. This term characterizes how the learning-rate and batch-size schedules shape the accumulation and dissipation of stochastic noise. The integrand $e(z)+\sigma^{2}]\gamma(z)$ represents the instantaneous noise magnitude, where $e(z)$ captures mini-batch noise and $\sigma^{2}$ captures label noise. The forgetting kernel $\mathcal{K}(\cdot)$ quantifies how noise injected at time $z$ still affects the loss at time $t$. Due to $\mathcal{K}(t)\asymp t^{-(2-1/\beta)}$, a higher-capacity model (smaller $\beta$) tends to forget noise more slowly.
Notably, the last two terms together constitute the optimization error and two key factors govern the trade-off between the them:
  • Model capacity. Increasing model capacity ($\beta\downarrow$) accelerates signal learning but simultaneously slows noise forgetting.
  • Learning-rate and batch-size schedules. Smaller learning rates or larger batch sizes suppress noise injection but also shorten the intrinsic training time. However, sufficient intrinsic time is important: the signal-learning term requires it to effectively reduce the risk, while the noise-forgetting term relies on it to forget noise memorized in early training. Hence, effective schedules must balance these competing objectives—suppressing injected noise while maintaining enough intrinsic time for both learning and forgetting.
4.2 General Results: The Top-$M$ Feature Case
The FSL (9) is established only for the hard-learning regime where $s\leqslant 1-1/\beta$. We now show that analogous FSLs also hold in general cases. To state the result, we define
$e_{M}(t)=\sum_{j=1}^{M}\lambda_{j}|\theta_{j}^{*}|^{2}e^{-2\lambda_{j}t},\qquad\mathcal{K}_{M}(t)=\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}t}.$
One can verify that both functions exhibit power-law decay for $1\lesssim t\lesssim M^{\beta}$ (see the right figure and Lemma C.4):
$e_{M}(t)\,\asymp\,t^{-s},\qquad\mathcal{K}_{M}(t)\,\asymp\,t^{-(2-1/\beta)},\qquad 1\lesssim t\lesssim M^{\beta}.$ (10)
Consequently, $e_{\infty}(t)\eqsim e(t)$ and $\mathcal{K}_{\infty}(t)\eqsim\mathcal{K}(t)$ for $t\geqslant 0$.
The following theorem provides a characterization of the loss dynamics for general case:
Theorem 4.3 (Intrinsic-Time FSL, top-$M$ features, general label noise).
[…]
The above characterization is uniform with respect to the label-noise level $\sigma$, and holds for all $s>0$ and $\beta>1$. It asserts that the exact FSL relation (11) (i.e., the FSL (9)) remains valid up to the intrinsic time $t\leqslant cM^{\beta}=:t_{M}$. For later times $t>t_{M}$, although the FSL may no longer hold exactly, the loss dynamics remain controlled from both sides as given in (12).
At the critical time $t_{M}$, we have $e_{M}(t_{M})\!\asymp\!M^{-s\beta}$, indicating that signal learning has reached the approximation-error limit. Beyond this point, further training no longer improves the learned signal; instead, the dynamics become dominated by noise effects. Depending on the interaction between the stochastic noise and the decaying learning rate, additional training may either inject more noise or dissipate it. Thus, it is a priori unclear whether the total error will significantly increase or decrease after $t_{M}$; however, the upper bound in (12) guarantees that the overall loss remains well-controlled, analogous to the behavior of the infinite-width limit ($M\to\infty$).
Nevertheless, an FSL may still hold for all $t\geqslant 0$, under suitably stronger conditions. In Theorem 4.2, we considered the setting with tasks satisfying $s\leqslant 1-1/\beta$. The following result shows that a similar characterization extends to general task difficulty with constant label noise.
Theorem 4.4 (Intrinsic-Time FSL, top-$M$ features, constant label noise).
Under Assumption 4.1, suppose $\sigma\gtrsim 1$. Let $\bm{\nu}_{t}$ denote the solution to the intrinsic-time SDE (8) with the top-$M$ features. Then, for any $s>0$ and all $t\geqslant 0$,
$\mathbb{E}[\mathcal{R}(\bm{\nu}_{t})]-\tfrac{1}{2}\sigma^{2}\;\eqsim\;M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-z)\,[e_{M}(z)+\sigma^{2}]\,\gamma(z)\,\mathop{}\!\mathrm{d}z.$
Theorem 4.2 implies that the finite-$M$ functions $e_{M}$ and $\mathcal{K}_{M}$ can be replaced by their infinite-width counterparts $e_{\infty}$ and $\mathcal{K}_{\infty}$ in the hard-learning regime. The next result demonstrates that the same FSL characterization naturally extends to the noiseless case $\sigma=0$.
Theorem 4.5 (Intrinsic-Time FSL, top-$M$ features, zero label noise).
Suppose Assumption 4.1 holds and $\sigma=0$. Let $\bm{\nu}_{t}$ denote the solution to the intrinsic-time SDE (8) with the top-$M$ features. If $s\in[0,2-1/\beta]$, then for all $t\geqslant 0$,
$\mathbb{E}[\mathcal{R}(\bm{\nu}_{t})]\;\eqsim\;M^{-s\beta}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-z)\,e_{M}(z)\,\gamma(z)\,\mathop{}\!\mathrm{d}z.$
4.3 General Results: The Random-$M$ Feature Case
For the random-features case, the modified feature covariance matrix is $\widehat{\mathbf{H}}=\mathbf{W}\mathbf{H}\mathbf{W}^{\top}$, whose eigenvalues we denote by $\widehat{\lambda}_{1}\geqslant\widehat{\lambda}_{2}\geqslant\cdots\geqslant\widehat{\lambda}_{M}$. We similarly define:
$\widehat{e}_{M}(t)=\sum_{j=1}^{M}\widehat{\lambda}_{j}|\theta_{j}^{*}|^{2}e^{-2\widehat{\lambda}_{j}t},\qquad\widehat{\mathcal{K}}_{M}(t)=\sum_{j=1}^{M}\widehat{\lambda}_{j}^{2}e^{-2\widehat{\lambda}_{j}t}.$
The next theorem establishes that the same FSL characterization also holds when the top-$M$ features are replaced by randomly selected features.
Theorem 4.6 (Intrinsic-Time FSL, random-$M$ features).
Suppose Assumption 4.1 holds and $s\in(0,1]$. Let $\bm{\nu}_{t}$ denote the solution to the intrinsic-time SDE (8) with the random-$M$ features. Then, with probability at least $1-\exp(-\Omega(M))$ over the randomness of the projection matrix $\mathbf{W}$, the results of Theorems 4.3, 4.4, and 4.5 continue to hold, after replacing $e_{M}(\cdot)$ and $\mathcal{K}_{M}(\cdot)$ with their random-feature counterparts $\widehat{e}_{M}(\cdot)$ and $\widehat{\mathcal{K}}_{M}(\cdot)$, respectively.
Lemma 4.7.
With probability at least $1-\exp(-\Omega(M))$ over the randomness of the projection matrix $\mathbf{W}$, it holds that $\widehat{\lambda}_{j}\eqsim\lambda_{j}\eqsim j^{-\beta}$ for any $j\in[M]$.
Theorem 4.6 and Lemma 4.7 together imply that when the task difficulty satisfies $s\leqslant 1$, training with random-$M$ features is similar to using the top-$M$ features, up to exponentially small probability. We emphasize, however, that for easier tasks with $s>1$, the behaviors of random and top feature may diverge and we leave this for future investigation.
4.4 Key Proof Steps and Core Insights
In this section, we outline the main ideas behind the proof of the FSL (11), highlighting the key techniques. Complete proofs of the above theorems are deferred to Appendix C.
We will need the following characterization of the gradient noise structure.